OpenStudy (anonymous):

How do I solve? 7x/2 + 6 = 1/4 by 4

OpenStudy (anonymous):

7X/2+6=1/4x4 4(7X/2+6/1=4/4) 14X+24=4 14X=-20 14X/14=-20/14 X=-1 and 3/10